$SPAD/src/input fib.input

Timothy Daly December 3, 2016

Abstract

Contents

2

1 License

— license —

--Copyright The Numerical Algorithms Group Limited 1991.

———-

— * —

)set break resume
)spool fib.output
)set message test on
)set message auto off
)clear all
--S 1 of 4
fib(n | n=0)==1
--R
--R                                                                   Type: Void
--E 1
--S 2 of 4
fib(n | n=1)==1
--R
--R                                                                   Type: Void
--E 2
--S 3 of 4
fib(n | n>1)==fib(n-1)+fib(n-2)
--R
--R                                                                   Type: Void
--E 3
--S 4 of 4
fibs == [fib i for i in 0..]
--R
--R                                                                   Type: Void
--E 4
)spool
)lisp (bye)

———-

References^{[1] nothing}

4